explain the processof carbon dating. c+* has a half-life of 5730 years? an archaeological sample is found with a carbon 14 to carbon12 ratio of 1/8. what is the age of the sample?


1
Expert's answer
2021-06-18T11:28:04-0400
"\frac{1}{8}=\frac{1}{2^3}\\t=3\tau=3(5730)=17190\ years"